# Manolis Antonoyiannakis American Physical Society, 1 Physics Ellipse, College Park, MD 20740

### v1

@@ 1.1 Impact Factors are scale-dependent
[…]
@@
What can be done to remedy this situation and level the playing field? One approach would be to abandon citation averages (and thus Impact Factors) altogether in assessing research journals. Indeed, several other flaws of IF rankings have long been identified, such as the skewness of citation distributions, the subject dependence of citation averages, the numerator/denominator asymmetry, journal self-citations, the length of citation window, IF inflation, IF gaming and engineering, or the document-type citation disparity (Amin & Mabe, 2004; Moed, 2005; Larivière & Sugimoto, 2019; Pulverer, 2013; Garfield, 1996; Gingras, 2016; Wouters et al., 2019). Growing concern for these shortcomings and their repercussions in IF rankings have led to public calls for abandoning, de-emphasizing, or at least complementing IFs with more reliable metrics (San Francisco Declaration on Research Assessment, 2012; Gaind, 2018) and detailed, statistically sensible visualizations (Adams et al., 2019).
@@ 2.1 What is a subject?
3 Unlike the Impact Factor, the $\Phi$ index is scale independent and passes the ‘random journal test’
@@
Does the $\Phi$ index pass this ‘random journal test’? If so, then plotting the $\Phi$ index vs. size would yield no size dependence. In Figures 5, 6, 7, and 8, we plot the citation average $f(n)$ (left panels) and $\Phi$ index (right panels) vs. journal size for 20,000 ‘journals’ drawn randomly from papers in the subjects of multidisciplinary sciences, condensed matter physics, psychology, and information science & library science, respectively. Evidently, the range of values of the $\Phi$ index in all these cases shows no dependence with journal size, unlike the citation average. Furthermore, the vast majority of $\Phi$ index values are bounded within the range $[-3,3]$. These results make sense and confirm our expectation from Eq. (6), where the inequality covers 99% of random journals for $k=3$.
@@ 6 Conclusions and outlook
In this work, we introduced the $\Phi$ index, a scale- and subject-independent citation average indicator. We explained how the $\Phi$ index is based on the Central Limit Theorem of Statistics, via standardization of citation averages in a process that is similar to calculating the $z$ score and that has an elegant geometric interpretation. We described a simple numerical check for testing any citation indicator for scale- and subject-independence. We then applied the $\Phi$ index to scientific journals. We compared $\Phi$-index with citation-average ($f$) rankings for journals using data from the 2020 Journal Citation Reports. We showed how the scale standardization of the $\Phi$ index corrects for the scale bias in $f$ rankings, and how the subject standardization allows journals from a more diverse range of subjects (such as mathematics, sociology, entomology, family studies, agronomy, history, linguistics, and law) in the top 1000 slots of $\Phi$ rankings. The $\Phi$-index methodology is statistically sensible, easy to implement, and can be extended to other populations such as researchers, universities, or countries.
Future work could involve the application of the $\Phi$ index to individual researchers or universities, or the comparison of the $\Phi$ index with other subject-normalized indicators for journals. It would also be worthwhile to study the extent to which the $\Phi$ index is indicative of the “high-end” of the underlying citation distributions, i.e., of papers that are highly cited, or even highlighted by the reputable science press. A preliminary study (Antonoyiannakis, 2023) showed that the $\Phi$ index is also a better descriptor of a journal’s potential for capturing highly cited papers (at the 99th percentile in their academic field per publication year) than the IF.

### v2

@@ 1.1 Impact Factors are scale-dependent
Every year at the end of June, when the Journal Impact Factors are announced by Clarivate Analytics, a familiar pattern occurs (Antonoyiannakis, 2018): (a) The high-Impact-Factor ranks are always populated by small journals that typically publish no more than a few hundred papers per discipline per year. For example, all the top 50 IF ranks—and all but 13 of the top 51–250 ranks—of the 2020 Journal Citation Reports (JCR) are filled by journals that publish fewer than 1000 papers/year. Of the 13 slots in the ranks 51–250 that are not occupied by small journals, we find 12 mid-sized journals (between 1000-5000 papers/year), and only one large journal ($>5000$ papers/year), in the 196th rank. (b) Small journals also occupy the low-Impact-Factor ranks—for example, all but one of the bottom 1000 ranks in the 2020 JCR, the one exception being a mid-sized journal. (c) Large journals that publish more than 5000 papers/year have modest Impact Factors that are neither very small nor very large, and tend to occupy unremarkable positions in the ranks—for example, there are 10 large journals with IF ranging from 2.5–14 and rank positions from 196 to 5189. Thus, large journals are precluded from reaching high IF ranks but are also “protected” from falling to the lowest of ranks as well. (d) Single papers affect the Impact Factors of hundreds—sometimes thousands—of small journals. For example, in 2017, the IF of 818 journals increased by more than 25% by their single top-cited paper, while one in 10 journals had their IF boosted by more than 50% by their top three cited papers (Antonoyiannakis, 2020).
@@
What can be done to remedy this situation and level the playing field? One approach would be to abandon citation averages (and thus Impact Factors) altogether in assessing research journals. Indeed, several other flaws of IF rankings have long been identified, such as the skewness of citation distributions, the subject dependence of citation averages, the numerator/denominator asymmetry, journal self-citations, the length of citation window, IF inflation, IF gaming and engineering, or the document-type citation disparity (Amin & Mabe, 2004; Moed, 2005; Larivière & Sugimoto, 2019; Pulverer, 2013; Garfield, 1996; Gingras, 2016; Wouters et al., 2019; Siler & Larivière, 2022). Growing concern for these shortcomings and their repercussions in IF rankings have led to public calls for abandoning, de-emphasizing, or at least complementing IFs with more reliable metrics (San Francisco Declaration on Research Assessment, 2012; Gaind, 2018) and detailed, statistically sensible visualizations (Adams et al., 2019).
@@ 2.1 What is a subject?
3 Unlike the Impact Factor, the $\Phi$ index is scale independent and passes the ‘random sample test’
@@
Does the $\Phi$ index pass this ‘random sample test’? If so, then plotting the $\Phi$ index vs. size would yield no size dependence. In Figures 5, 6, 7, and 8, we plot the citation average $f(n)$ (left panels) and $\Phi$ index (right panels) vs. journal size for 20,000 ‘journals’ drawn randomly from papers in the subjects of multidisciplinary sciences, condensed matter physics, psychology, and information science & library science, respectively. Evidently, the range of values of the $\Phi$ index in all these cases shows no dependence with journal size, unlike the citation average. Furthermore, the vast majority of $\Phi$ index values are bounded within the range $[-3,3]$. These results make sense and confirm our expectation from Eq. (6), where the inequality covers 99% of random journals for $k=3$.
@@ 6 Conclusions and outlook
In this work, we introduced the $\Phi$ index, a scale- and subject-independent citation average indicator. We explained how the $\Phi$ index is based on the Central Limit Theorem of Statistics, via standardization of citation averages in a process that is similar to calculating the $z$ score and that has an elegant geometric interpretation. Thus, the $\Phi$ index of a sample of $n$ scientific papers is the distance of the sample’s citation average $f$ from the subject citation mean $\mu$, measured (normalized) in units of the sample’s standard deviation $\sigma_{n}=\sigma/\sqrt{n}$.
We described a simple numerical check for testing any citation indicator for scale- and subject-independence. We then applied the $\Phi$ index to scientific journals. We compared $\Phi$-index with citation-average ($f$) rankings for journals using data from the 2020 Journal Citation Reports. We showed how the scale standardization of the $\Phi$ index corrects for the scale bias in $f$ rankings, and how the subject standardization allows journals from a more diverse range of subjects (such as mathematics, sociology, entomology, family studies, agronomy, history, linguistics, and law) in the top 1000 slots of $\Phi$ rankings. The $\Phi$-index methodology is mathematically sound, statistically robust, easy to implement, and can be extended to other populations such as researchers, universities, or countries.
Future work could involve the application of the $\Phi$ index to individual researchers or universities, or the comparison of the $\Phi$ index with other subject-normalized indicators for journals. It would also be interesting to study whether and to what extent the $\Phi$ index is vulnerable to gaming. Another project would be to study the extent to which the $\Phi$ index is indicative of the “high-end” of the underlying citation distributions, i.e., of papers that are highly cited, or even highlighted by the reputable science press. A preliminary study (Antonoyiannakis, 2023) showed that the $\Phi$ index is also a better descriptor of a journal’s potential for capturing highly cited papers (at the 99th percentile in their academic field per publication year) than the IF.